silemus.

30Uniform probability P(X>7)P(X > 7)

X∼Uniform(0,10)\displaystyle X \sim \mathrm{Uniform}(0, 10). Find P(X>7)\displaystyle P(X > 7).

For a Uniform(0,10), probability of any interval is its length over the total length. The region X > 7 has length 10 − 7 = 3, so P(X > 7) = 3/10 = 0.3. Check: P(X ≤ 7) + P(X > 7) = 0.7 + 0.3 = 1.

How good is this answer?